### v1

@@ 2.2 RL with Unsupervised Reward for Key Slice Summary
where $R_{sim}$ calculates the cosine similarities ($sim$) between features to constrain redundancy, $R_{cls}$ uses the classification loss to feedback anatomical retention ($C$ and $y_{c}$ represents the category and the corresponding probabilities). Specifically, by taking the maximum value at the time dimension, the features of key slices of size ($\mathcal{S},\mathcal{D}$) will be fused to $\hat{f}_{\mathcal{S}}=(1,\mathcal{D})$, which then be combined with the volume features and inputted into a linear layer with Softmax for classification output (named fusion linear module, FLM). $R_{penalty}$ penalizes the situations of storing over $\mathcal{S}_{max}=5$ slices. Since no key slice labels are available, the reward functions are conducted in an unsupervised manner. Our goal is to learn a policy ($\pi_{\zeta}$) that maximizes the expected reward $J(\zeta)$. We follow REINFORCE algorithm [23] and approximate the gradient by running $N$=100 episodes and averaging the results:
@@ 2.3 Uncertainty-aware Strategy for Classification Adjustment
$S_{unc}(c_{t_{i}})=\sum_{i=1}^{K}\left\|p_{k}(c_{t_{i}})-\frac{1}{K}\sum_{i=1}^{K}p_{k}(c_{t_{i}})\right\|^{2},$ (4)
@@ 3 Experimental Results
Materials and Implementation Details. Our private large 3D uterus dataset in pelvic US contains 677 volumes, annotated with coronal standard planes, and uterine categories: normal (N, 345), unicornuate (U, 52), bicornuate (B, 5), arcuate (A, 145), T-shaped (T, 15), septate uterus (S, 96), and uterus didelphys (D, 19), by experienced sonographers under strict quality control using the Pair annotation software package [18]. The average volume size is 367$\times$189$\times$334, with spacing equal to 0.43 $mm^{3}$. We randomly split the volumes into 406, 68, and 203 for training, validation, and testing, respectively.
@@
In this study, we implemented our method in PyTorch, using one NVIDIA 4090 GPU with 24GB memory. In stage 1, the denoising diffusion model with cosine noise scheduler was developed based on diffusers package by HuggingFace. The time steps are set to 1000 for training, and 100 for validation and testing. The training includes 200K interactions with batch size=8 and learning rate ($lr$)=5e-4. The architectures of parameter/plane/volume encoders follow the design in [7], here, we used different multi-layer perceptions (MLPs) to map them to $\mathbb{R}^{32}$. The text encoder followed BiomedCLIP [28], obtaining $\mathbb{R}^{32}$ features via MLP. Through concatenation, they were input into the UNet-based denoising network (with the positional encoding of timestep added to all blocks). In stage 2, we trained the RL agent and FLM with lr=5e-4 in 100K interactions. Alternate training was conducted every 1K iterations. Both stages used AdamW optimizers and conducted validation every 2K iterations, and models with the best validation performance were saved for final evaluation. All slices were resized to 2242, and volume size standardization was not required since we processed one volume per iteration. They were then scaled to 0–1 by dividing 255.
@@
Table 3: Different localization methods with ($w$) and without ($w/o$) slice summary (SS) for classification. The improved metrics are highlighted in bold.
@@
Table 1 compares our method with five iterative approaches, i.e., one traditional CNN [17], three RL methods [25, 26, 29], and one diffusion-based solution [7]. Results show that our proposed method significantly outperforms all the competitors ($p<0.05$) in terms of spatial and content similarities. In Table 2, we evaluated various typical and advanced methods and their variants, containing three 2D/2.5D traditional networks [10], three video-based approaches [2, 19, 16], and seven 3D models [10, 4, 11, 6, 9]. Results show that they all obtain Acc and F1 less than 90%. Our proposed method shows significant improvement over them in all evaluation metrics (last row, $p<0.05$), validating its effectiveness. Table 2 also provides the ablation studies to test the contribution of each proposed strategy. $AP$, $PP$, and $SS$ represent using all planes, the final predicted plane, and slice summary as classification inputs, respectively. $GF$ means the global volumetric features, and $UA$ denotes the uncertainty-aware strategy. We observe that even with $GF$, using only one plane degrades the performance compared to using all planes. Besides, $SS$ can remove abundant information, extract key features, and improve performance (F1: 22.45$\uparrow$). Notably, with $UA$ only, our model achieves satisfactory performance with Acc, F1 and Pre$>$90%. Adding $SS$ and $GF$ can further enhance the performance on all evaluation metrics. We further prove that $SS$ is general to different plane localization methods, with significantly improved performance (Table 3, all p-values$<$0.05).

### v2

@@ 2.2 RL with Unsupervised Reward for Key Slice Summary
where $R_{sim}$ calculates the cosine similarities (simsim) between features to constrain redundancy, $R_{cls}$ uses the classification loss to feedback anatomical retention ($C$ and $y_{c}$ represents the category and the corresponding probabilities). Specifically, by taking the maximum value at the time dimension, the features of key slices of size ($\mathcal{S},\mathcal{D}$) will be fused to $\hat{f}_{\mathcal{S}}=(1,\mathcal{D})$, which then be combined with the volume features and inputted into a linear layer with Softmax for classification output (named fusion linear module, FLM). $R_{penalty}$ penalizes the situations of storing over $\mathcal{S}_{max}=5$ slices. Since no key slice labels are available, the reward functions are conducted in an unsupervised manner. Our goal is to learn a policy ($\pi_{\zeta}$) that maximizes the expected reward $J(\zeta)$. We follow REINFORCE algorithm [23] and approximate the gradient by running NN=100 episodes and averaging the results:
@@ 2.3 Uncertainty-aware Strategy for Classification Adjustment
$S_{unc}(c_{t_{i}})=\sum_{i=1}^{K}\left\|p_{k}(c_{t_{i}})-\frac{1}{K}\sum_{k=1}^{K}p_{k}(c_{t_{i}})\right\|^{2},$ (4)
@@ 3 Experimental Results
[…]
@@
In this study, we implemented our method in PyTorch, using one NVIDIA 4090 GPU with 24GB memory. In stage 1, the denoising diffusion model with cosine noise scheduler was developed based on diffusers package by HuggingFace. The time steps are set to 1000 for training, and 100 for validation and testing. The training includes 200K interactions with batch size=8 and learning rate (lrlr)=5e-4. The architectures of parameter/plane/volume encoders follow the design in [7], here, we used different multi-layer perceptions (MLPs) to map them to $\mathbb{R}^{32}$. The text encoder followed BiomedCLIP [28], obtaining $\mathbb{R}^{32}$ features via MLP. Through concatenation, they were input into the UNet-based denoising network (with the positional encoding of timestep added to all blocks). In stage 2, we trained the RL agent and FLM with lr=5e-4 in 100K interactions. Alternate training was conducted every 1K iterations. Both stages used AdamW optimizers and conducted validation every 2K iterations, and models with the best validation performance were saved for final evaluation. All slices were resized to 2242, and volume size standardization was not required since we processed one volume per iteration. They were then scaled to 0–1 by dividing 255.
@@
Table 3: Different localization methods with ($w$) and without ($w/o$) slice summary ($S$) for classification. The improved metrics are highlighted in bold.
@@
[…]